The mean weight of the luggage carried into an airplane by individual passengers at Taguegarao City Airport is 19.8 kilograms. A statistician takes a random sample of 110 passengers and obtain a sample mean weight of 18.5 kilograms with standard deviation of 8.5 kilograms. Test the claim at a=0.01 level of significance.

1
Expert's answer
2022-05-18T08:55:07-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=19.8"

"H_a:\mu\not=19.8"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=109" degrees of freedom, and the critical value for a two-tailed test is "t_c =2.621688."The rejection region for this two-tailed test is "R = \{t:|t|> 2.621688\}."

The t-statistic is computed as follows:


"t=\dfrac{18.5-19.8}{8.5/\sqrt{110}}=-1.6041"

Since it is observed that "|t| = 1.6041<2.621688=t_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for two-tailed "df=109" degrees of freedom, "t=-1.6041" is "p=0.111586," and since "p=0.111586>0.01=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu" is different than 19.8, at the "\alpha = 0.01" significance level.